OpenStudy (anonymous):

If the area (in square units) of the region under the curve of the function f(x) = 4 on the interval [1, a] is 20 square units, and a > 1, then what is the value of a

OpenStudy (anonymous):

a=6 from Area = y * x (length) = 4 * (a-1) a= 6